# Graph-Based Bidirectional Transformer Decision Threshold Adjustment Algorithm for Class-Imbalanced Molecular Data

### v1

@@ 3.2 Evaluation Metrics
There are various existing evaluation metrics for data classification tasks. One such metric is model accuracy, which computes the fraction of data points correctly classified by the model. However, this metric does not capture many nuances of model performance. Particularly when dealing with imbalanced data, as Esposito et al. discussed in [12], the minority class may be the class of most interest, but points from the minority class are more likely to be misclassified by the model than points from the majority class(es). In highly imbalanced data sets, such as the ones used in the present work, there are comparatively very few points in the minority class. Even if the model misclassifies all of the points in the minority class, the accuracy of the model could still be very high, and the model could mistakenly appear to be suitable for the highly imbalanced classification task. Thus, to quantify our model performance in the present work, we use another evaluation metric: the area under the receiver operating characteristic (ROC) curve, also known as the ROC-AUC (area under curve) score.
@@
For a classification task, the ROC curve plots the true positive rate (TPR) versus the false positive rate (FPR) of the classifier for various thresholds. The ROC curve captures useful information about the classifier, particularly on highly imbalanced data–because the minority (or positive, in our case) class can be easily misclassified, both the TPR and FPR are important to consider. Recall that a classification threshold is used to predict a given data point’s class using its probability distribution over all classes. If the probability that the data point belongs to the minority class is greater than the threshold, then that data point is assigned to the minority class. For a threshold equal to 0, the model would predict that every data point belonged to the minority class. In this case, the model would successfully classify all of the true positive data points (i.e., a TPR of 1), but it would also incorrectly classify all of the negative data points as positive (i.e., an FPR of 1). As the threshold increased, fewer data points would be classified as positive. For a threshold equal to 1, the model would classify all data points as negative, producing a TPR and an FPR of 0.
@@
The ROC-AUC score quantifies the relationship between the TPR and FPR by computing the area under the ROC curve. For an area close to 1, it is theoretically possible to choose an ”optimal” threshold that yields a high TPR and a low FPR (i.e., produces a point close to the upper-left corner of the ROC plot). In general, an AUC of 0.5 suggests that that algorithm has no ability to distinguish between the classes, 0.7 to 0.8 indicates an acceptable model, 0.8 to 0.9 describes an excellent model, and more than 0.9 designates an outstanding model. To calculate the ROC-AUC score in our experiments for the proposed BTDT-MBO method, we used the perfcurve function in MATLAB. There also exist functions that calculate the ROC-AUC score for the multiclass case, so our evaluation procedure can be applied for any number of classes.
@@
To further illustrate the construction of ROC curves and their reflection of the classification performance of models, we have included example ROC curves in Figure 1 for two of the data sets used in benchmarking the proposed BTDT-MBO model. These plots were generated using a random labeled/unlabeled split of each data set, and the thresholds used to construct each plot ranged from 0 to 1, with each threshold increasing by 0.05. The ROC curve on the right for the CHEMBL1909150 data set demonstrates near-ideal performance on this data partition, with a point very close to the upper-left corner of the plot, corresponding to an FPR close to 0 and a TPR close to 1. Indeed, as displayed in Figure 2 and discussed in Section 3.3, our model’s average ROC-AUC score for 50 random partitions of the CHEMBL1909150 data set was 0.978, displaying outstanding classification performance. The ROC curve on the left for the CHEMBL1909134 data set clearly has a lower ROC-AUC value than the curve for the CHEMBL1909150 data set, but the model still shows some ability to discriminate between the two classes. As shown in Figure 2, our model’s average ROC-AUC score over all partitions on the CHEMBL1909134 data set was 0.792, which is not surprising given its example ROC curve. Do note, however, that our proposed model performed better than the comparison methods on this data set, which is further discussed in Section 3.3.
3.3 Performance and Discussion
@@
To evaluate the performance of our proposed BTDT-MBO method for highly imbalanced molecular data sets, we compare our models against the GHOST algorithm from Esposito et al. [12]. The GHOST algorithm is designed to be paired with any machine learning classifier, and the code used in [12] ran experiments using GHOST with logistic regression (LR), random forest (RF), gradient boosting (GB), and extreme gradient boosting (XGB) classifiers. In the present work, all models used 80% of the data as training (or labeled, in the MBO case) to mirror the experiments by Esposito et al. [12].
@@
Figure 2: Comparison of our proposed method (shown in red) with other methods (shown in blue) on four DrugMatrix data sets. The imbalance ratios for the pictured data sets vary from 16.5 to 20.0. Detailed information about the size and composition of the data sets is given in Section 3.1. Performance is given as the ROC-AUC score averaged over 50 random training-testing (or labeled/unlabeled) splits of the given data set, with 80% of the data being labeled in each case. The BTDT-MBO result for each data set is the highest of the BTDT-MBO model using a Gaussian weight function and the BTDT-MBO model using a distance correlation weight function. The comparison results were generated using the GHOST algorithm [12] with logistic regression (LR), random forest (RF), gradient boosting (GB), and extreme gradient boosting (XGB).
@@
Figure 3: Comparison of our proposed method (shown in red) with other methods (shown in blue) on two DS1 data sets. The imbalance ratios for the pictured data sets are both 20.0. Detailed information about the size and composition of the data sets is given in Section 3.1. Performance is given as the ROC-AUC score averaged over 50 random training-testing (or labeled/unlabeled) splits of the given data set, with 80% of the data being labeled in each case. The BTDT-MBO result for each data set is the highest of the BTDT-MBO model using a Gaussian weight function and the BTDT-MBO model using a distance correlation weight function. The comparison results were generated using the GHOST algorithm [12] with logistic regression (LR), random forest (RF), gradient boosting (GB), and extreme gradient boosting (XGB).
@@
Figures 2 and 3 display the results of the proposed BTDT-MBO method compared with the GHOST method from Esposito et al. [12]. We generated results using GHOST paired with LR, RF, GB, and XGB. The GHOST model allows the user to select the desired classification metric to use for optimization from two options: the Cohen’s kappa, which is a balanced classification metric that was used as the default metric in [12], or the ROC-AUC score. All of the experiments in the present work were evaluated using the ROC-AUC score.
@@
Overall, the results show that the proposed BTDT-MBO method obtains higher ROC-AUC scores than the comparison methods for all data sets, except the CHEMBL8 data set, for which the result of the proposed algorithm is second best, but almost the same as the best. We will first outline the details of the experiments in the present work, after which we will provide a summary and further discussion of the results.
@@
To create labeled and unlabeled partitions in the data input to our DT-MBO method, we randomly selected 80% of the points to be labeled and denoted the rest as unlabeled. Then, we applied the DT-MBO method, which propagates the known labels to the unlabeled nodes for a prescribed number of iterations as described in Section 2.1.1. We used the resulting predicted labels to calculate the optimal ROC-AUC score for this partition (over all thresholds). In highly imbalanced data sets, which often contain very few points in the minority class, results can vary widely depending on the particular training/testing or labeled/unlabeled partition. To account for this variability, as in [12], we repeated each experiment 50 times, with a new random partition for each experiment, and averaged the ROC-AUC scores over the 50 trials. We used a similar splitting and averaging procedure to generate the GHOST comparison results; however, Esposito et al. [12] recorded the random seeds used for each training/testing data split in their experiments, and we used the same 50 random seeds when running the GHOST tests. Results in Figures 2 and 3 are given as the average ROC-AUC score over 50 data splits.
@@
Regarding parameters, the proposed BTDT-MBO model requires specification of various parameters, whose functions can be seen in Algorithms 2 and 1. The bidirectional transformer [8] requires the user to choose one of three pretraining data sets as outlined in Section 2.1.2. In the present work, all BTDT-MBO experiments use the transformer model trained only on the ChEMBL data set [24], based on its success in predicting scarcely labeled molecular data in experiments from the authors’ previous work [28] using the BT-MBO method. The transformer model does not require any additional parameter specifications, as all other parameters are implicit in the pretrained models [8, 21].
@@
The DT-MBO algorithm relies on several parameters, many of which are discussed in our previous work [28], but some of which have been added in the present work to implement the decision threshold adjustment. Furthermore, some parameters were defined manually prior to running all experiments, and other parameters were hyperparameters that were tuned through our experiments. Here, we provide a brief description of all parameters and their settings or tuning. First, to create the framework in which the DT-MBO algorithm operates, we construct a graph with $N_{n}$ neighbors for each data point. Then, the graph Laplacian $\mathbf{L}$ (2) is constructed, and the first (smallest) $N_{e}$ eigenvalues and corresponding eigenvectors of $\mathbf{L}$ are computed. Overall, both $N_{n}$ and $N_{e}$ were designated as hyperparameters and tuned in our experiments. Other hyperparameters for the DT-MBO method that were tuned in our experiments are $C$ (the factor in the diffusion step), $dt$ (the time step used in applying the diffusion step), and $N_{t}$ (the number of iterations of the diffusion and thresholding steps the model should repeat).
@@
Parameters that were set manually include the number of times $N_{s}$ to apply the diffusion operator in the BTDT-MBO method, which was set at 3, and the number of labeled sets $N_{l}$, which was set to 50 for all tasks. Additionally, the number of labeled points $N_{p}$ was set at 80% of each data set for all tasks. Finally, the decision threshold adjustment steps in our algorithm require the specification of a minimum threshold and a maximum threshold to test. For all BTDT-MBO experiments, we used a minimum threshold of 0.05 and a maximum threshold of 0.55.
@@
As previously stated, the proposed BTDT-MBO algorithm obtained higher ROC-AUC scores than the compared methods in all but one of the six highly imbalanced data sets included in this work. Figure 2 displays our model’s results (using the best-performing of the BTDT-MBO with Gaussian weight function and BTDT-MBO with distance correlation weight function) compared with the GHOST algorithm’s results on four highly imbalanced DrugMatrix data sets. The BTDT-MBO method achieved a higher ROC-AUC score than all GHOST models for all four of these DrugMatrix data sets. As shown in Table 1, The four DrugMatrix data sets each contained 842 compounds and had IRs of at least 16.5. Our model achieved its best predictive performance on the CHEMBL1909150 data set with an average ROC-AUC score of 0.978, indicating an outstanding ability to discriminate between the active and inactive classes. Furthermore, our model outperformed all GHOST models, which also demonstrated very high ROC-AUC scores on this data set. For the CHEMBL1909157 data set, our model achieved excellent results, with an ROC-AUC score of 0.866. For this data set, three of the GHOST models also achieved ROC-AUC scores greater than 0.8, but the GHOST model using GB performed significantly worse. Our model similarly outperformed all four GHOST comparison models on the CHEMBL1909132 and CHEMBL1909134 data sets.
@@
Figure 3 shows our model’s results on two highly imbalanced DS1 data sets, again compared with the GHOST algorithm paired with LR, RF, XGB, and GB. Recall from Section 3.1 that all DS1 data sets have an IR of 20.0 by construction. For the CHEMBL100 data set, our model achieved an outstanding ROC-AUC score of 0.974, while the highest-performing GHOST model (using LR) earned an ROC-AUC score of 0.959. While our BTDT-MBO method did not achieve the highest ROC-AUC score of all tested algorithms on the CHEMBL8 data set, it had the second-best ROC-AUC score, performing almost as well as the top-performing GHOST model using LR. All models in this case demonstrated ROC-AUC scores of at least 0.9, with the highest method earning a score of 0.979 and BTDT-MBO scoring 0.978.
@@
As outlined in Section 2.2, in addition to the default Gaussian kernel used in our BTDT-MBO method, we also conducted experiments that instead used the distance correlation to compute the weights between data elements. In our experiments, the BTDO-MBO method using a Gaussian weight function yielded higher ROC-AUC scores than the BTDT-MBO method using distance correlation for five out of the six data sets included in this work, with the BTDT-MBO algorithm using distance correlation achieving the best results on the CHEMBL1909157 DrugMatrix data set. However, the BTDT-MBO algorithm using a distance correlation weight function performed similarly to (and frequently almost as well as) the BTDT-MBO method using a Gaussian kernel on all six data sets, particularly on the DS1 data sets used for benchmarking. For the CHEMBL100 DS1 data set, the BTDT-MBO model using distance correlation yielded an ROC-AUC score of 0.970, slightly lower than BTDT-MBO using Gaussian weights (ROC-AUC = 0.974) but still higher than all comparison methods. Similarly, for the CHEMBL8 DS1 data set, BTDT-MBO using distance correlation scored 0.973, which was again lower than BTDT-MBO using a Gaussian weight function (ROC-AUC = 0.978) but higher than two comparison methods. Again, all tested models scored very highly on this data set.
@@
Our model using distance correlation saw similar results on the DrugMatrix data sets. In fact, as mentioned above, BTDT-MBO using distance correlation achieved the best overall result of all tested models on the CHEMBL1909157 DrugMatrix data set, earning an ROC-AUC score of 0.878. BTDT-MBO using a Gaussian kernel scored an ROC-AUC value of 0.866 for this data set, so both BTDT-MBO methods beat all tested GHOST models. For the CHEMBL1909150 data set, BTDT-MBO using distance correlation scored an ROC-AUC value of 0.975, higher than all comparison methods and only lower than our BTDT-MBO model using a Gaussian, which scored 0.978. On the CHEMBL1909132 DrugMatrix data set, BTDT-MBO using distance correlation earned an ROC-AUC score of 0.743, which places it below BTDT-MBO using Gaussian weights (ROC-AUC = 0.785) and GHOST using LR, but above the three other GHOST comparison methods. Our model using distance correlation saw similar relative results for the CHEMBL1909134 DrugMatrix data set, with an average ROC-AUC score of 0.766. The model earned the third-highest score of all methods tested for this data set, lower than our BTDT-MBO method using Gaussian weights (ROC-AUC = 0.792) and very close to GHOST paired with LR, which scored a 0.769. Overall, these results suggest that the distance correlation can be useful for quantifying similarity in molecular data sets, including those that are highly imbalanced. Furthermore, the distance correlation may be considered as a potential alternative weight function to the Gaussian function in similarity graph-based settings, particularly when the data is not close to a normal distribution.
@@
Another novel addition to the proposed BTDT-MBO method is its threshold optimization step, the details of which are given in Section 2.1.1. This step chose the decision threshold that yielded the highest ROC-AUC score for a given partition of a particular data set. Because the partitions were created randomly, the class distributions of the labeled and unlabeled portions of a data set could vary. Additionally, due to the high level of imbalance in the data sets, there may be very few data points from the minority class in the labeled or unlabeled set for a particular partition. Thus, it is not surprising that in our experiments, the optimal thresholds for the splits varied considerably. However, the optimal thresholds tended to be lower than the typical threshold of 0.5. For example, in one experiment on the CHEMBL1909157 data set (i.e., testing over 50 random labeled/unlabeled partitions), the lowest optimal threshold for the 50 random splits was 0.05, and the highest optimal threshold was 0.45. The median optimal threshold over all 50 splits was 0.15. These results reinforce the impact of adjusting the decision threshold on classifier performance for highly imbalanced data.

### v3

@@ 3.2 Evaluation Metrics
There are various existing evaluation metrics for data classification tasks. One such metric is model accuracy, which computes the fraction of data points correctly classified by the model. However, this metric does not capture many nuances of model performance. Particularly when dealing with imbalanced data, as Esposito et al. discussed in 1, the minority class may be the class of most interest, but points from the minority class are more likely to be misclassified by the model than points from the majority class(es). In highly imbalanced data sets, such as the ones used in the present work, there are comparatively very few points in the minority class. Even if the model misclassifies all of the points in the minority class, the performance of the technique could still be very high, and the model could mistakenly appear to be suitable for the highly imbalanced classification task. Thus, to better quantify the performance of our proposed method, we use another evaluation metric, which computes the area under the receiver operating characteristic curve (the ROC curve). This is also known as the ROC-AUC score.
@@
For the data classification problem, the ROC curve plots the true positive rate versus the false positive rate of the classifier for various thresholds. The ROC curve captures useful information about the classifier, particularly on highly imbalanced data–because the minority (or positive, in our case) class can be easily misclassified, both the true positive rate (TPR) and the false positive rate (FPR) need to be considered. Recall that a classification threshold is used to predict a given data point’s class using its probability distribution over all classes. If the probability that the data point is affiliated with the minority class is greater than the threshold, then that point is allotted to the minority class. For a threshold equal to 0, the model would predict that every data point belonged to the minority class. In this case, the model would successfully classify all of the true positive data points (i.e., a TPR of 1), but it would also incorrectly classify all of the negative data points as positive (i.e., an FPR of 1). As the threshold increased, fewer data points would be classified as positive. For a threshold equal to 1, the model would classify all data points as negative, producing a TPR and an FPR of 0.
@@
The ROC-AUC score quantifies the relationship between the TPR and FPR by computing the area under the ROC curve. For an area close to 1, it is theoretically possible to choose an ”optimal” threshold that yields a high TPR and a low FPR (i.e., produces a point close to the upper-left corner of the ROC plot). In general, an AUC of 0.5 suggests that that algorithm has no ability to distinguish between the classes, 0.7 to 0.8 indicates an acceptable technique, 0.8 to 0.9 describes an excellent technique, and greater than 0.9 designates an outstanding technique. To calculate the ROC-AUC score in our experiments for the proposed BTDT-MBO method, we used the perfcurve function in MATLAB. There also exist functions that calculate the ROC-AUC score for the multiclass case, so our evaluation procedure can be applied for any number of classes.
@@
To further illustrate the construction of the ROC curves as well as their reflection of the classification performance of models, we have included example ROC curves in Figure 2 for two data sets used in benchmarking the proposed BTDT-MBO model. These plots were generated using a random labeled/unlabeled split of each data set, and the thresholds used to construct each plot ranged from 0 to 1, with each threshold increasing by 0.05. The ROC curve on the right for the CHEMBL1909150 data set demonstrates near-ideal performance on this data partition, with a point very close to the upper-left corner of the plot, corresponding to an FPR close to 0 and a TPR close to 1. Indeed, as displayed in Figure 3 and discussed in Section 3.3, our model’s average ROC-AUC score for 50 random partitions of the CHEMBL1909150 data set was 0.978, displaying outstanding classification performance. The ROC curve on the left for the CHEMBL1909134 data set clearly has a lower ROC-AUC value than the curve for the CHEMBL1909150 data set, but the model still shows some ability to discriminate between the two classes. As shown in Figure 3, our model’s average ROC-AUC score over all partitions on the CHEMBL1909134 data set was 0.792, which is not surprising given its example ROC curve. Do note, however, that our proposed model performed superior to the comparison methods on this data set, which is further discussed in Section 3.3.
3.3 Model Performance and Discussion
@@
To evaluate the performance of our BTDT-MBO method for highly imbalanced molecular data sets, we compare our models against the GHOST algorithm from Esposito et al. 1. The GHOST algorithm is designed to be paired with any machine learning classifier, and the code used in 1 ran experiments using GHOST with random forest (RF), extreme gradient boosting (XGB), logistic regression (LR) and gradient boosting (GB) classifiers. In the present work, all models used 80% of the data as training (or labeled, in the MBO case) to mirror the experiments by Esposito et al. 1.
@@
Figure 3: Comparison to other techniques on DrugMatrix data sets. The results of our proposed method are in red, while those of other algorithms are in blue. The imbalance ratios for the pictured data sets vary from 16.5 to 20.0. Detailed information about the overall size and composition of the comparison data sets is described in Section 3.1. The performance metric is the ROC-AUC score averaged over 50 random training-testing (or labeled/unlabeled) splits of the given data, with 80% of the data being labeled in each case. The BTDT-MBO result for each data set is the highest of the BTDT-MBO model using a Gaussian weight function and the BTDT-MBO model using a distance correlation weight function. Some comparison results were generated using the GHOST algorithm 1 via random forests (RF), extreme gradient boosting (XGB), logistic regression (LR) and gradient boosting (GB). Additional comparison results were generated using the BT-MBO algorithm 40 as well as BT-GB, BT-RF, and BT-SVM models (consisting of the BT-FPs passed to gradient boosting, random forest, and support vector machine algorithms, respectively).
@@
Figure 4: Comparison to other techniques on two DS1 data sets. The results of our proposed method are in red, while those of other algorithms are in blue. The imbalance ratios for the pictured data sets are both 20.0. Detailed information about the overall size and composition of the comparison data sets is described in Section 3.1. The metric is the ROC-AUC score averaged over 50 splits of the data, with 80% of the data being labeled in each case. The BTDT-MBO result for each data set is the highest of the BTDT-MBO model using a Gaussian weight function and the BTDT-MBO model using a distance correlation weight function. Some comparison results were generated using the GHOST algorithm 1 via random forests (RF), extreme gradient boosting (XGB), logistic regression (LR) and gradient boosting (GB). Additional comparison results were generated using the BT-MBO algorithm 40 as well as BT-GB, BT-RF, and BT-SVM models (consisting of the BT-FPs passed to gradient boosting, random forest, and support vector machine algorithms, respectively).
@@
Figures 3 and 7 display the results of the BTDT-MBO procedure compared with the GHOST method from Esposito et al. 1. We generated results using GHOST paired with LR, RF, GB, and XGB. The GHOST model allows the user to select the desired classification metric to use for optimization from two options: the Cohen’s kappa, which is a balanced classification metric that was used as the default metric in 1, or the ROC-AUC score. All of the experiments in this paper were evaluated using the ROC-AUC score.
@@
Overall, the experimental results indicate that the BTDT-MBO algorithm obtains higher ROC-AUC scores than the comparison methods for all data, except CHEMBL8, for which the result of the proposed algorithm is second best, but almost the same as the best. We first outline the details of the experiments in this work, after which we provide a summary as well as a further discussion of results.
@@
To create labeled and unlabeled partitions in the data input to our DT-MBO method, we randomly selected 80% of the points to be labeled and denoted the rest as unlabeled. Then, we applied the DT-MBO method, which propagates the known labels to the unlabeled nodes for a prescribed number of iterations as detailed in Section 2.1.1. We used the resulting predicted labels to calculate the optimal ROC-AUC score for this partition (over all thresholds). In highly imbalanced data sets, which often contain very few points in the minority class, results can vary widely depending on the particular training/testing or labeled/unlabeled partition. To account for this variability, as in 1, we repeated each experiment 50 times, with a new random partition for each experiment, and averaged the ROC-AUC scores over the 50 trials. We used a similar splitting and averaging procedure to generate the GHOST comparison results; however, Esposito et al. 1 recorded the random seeds used for each training/testing data split in their experiments, and we used the same 50 random seeds when running the GHOST tests. Results in Figures 3 and 7 indicate the average ROC-AUC score over 50 different splits of the data.
@@
Regarding parameters, the proposed BTDT-MBO model requires specification of various parameters, whose functions can be seen in Algorithms 2 and 1. The bidirectional transformer 43 requires the user to choose one of three pretraining data sets as outlined in Section 2.1.2. In the present work, all BTDT-MBO experiments use the transformer model trained only on the ChEMBL data set 58, based on its success in predicting scarcely labeled molecular data in experiments from the authors’ previous work 40 using the BT-MBO method. The transformer model does not require any additional parameter specifications, as all other parameters are implicit in the pretrained models 43, 70.
@@
The DT-MBO algorithm relies on several parameters, many of which are discussed in our previous work 40, but some of which have been added in the present work to implement the decision threshold adjustment. Some parameters were defined manually prior to running all experiments, and other parameters were tuned in our experiments. Here, we provide a brief description of all parameters and their settings or tuning. First, to create the framework in which the DT-MBO algorithm operates, we construct a $N_{n}$-neighbor graph. After, the graph Laplacian $\mathbf{L}$ (2) is constructed, and the first (smallest) $N_{e}$ eigenvalues and corresponding eigenvectors of $\mathbf{L}$ are computed. Overall, both $N_{n}$ and $N_{e}$ were designated as hyperparameters and tuned in our experiments. Other hyperparameters for the DT-MBO method that were tuned are the $C$ and $dt$ parameters in the diffusion step, as well as the number of iterations, $N_{t}$.
@@
Regarding other parameters, we let $N_{s}=3$ and $N_{l}=50$. Additionally, the number of labeled points $N_{p}$ was set at 80% of each data set for all tasks. Finally, the decision threshold adjustment steps in our algorithm require the specification of a minimum threshold and a maximum threshold to test. For all BTDT-MBO experiments, we used a minimum threshold of 0.05 and a maximum threshold of 0.55.
@@
As previously stated, the proposed BTDT-MBO algorithm obtained higher ROC-AUC scores than the compared methods in all but one of the six highly imbalanced data sets included in this work. Figure 3 displays our model’s results (using the best-performing of the BTDT-MBO with Gaussian weight function and BTDT-MBO with distance correlation weight function) compared with the GHOST algorithm’s results on four highly imbalanced DrugMatrix data sets. The BTDT-MBO procedure obtained a higher ROC-AUC score than all GHOST models for all four of the DrugMatrix data sets. As indicated in Table 1, each of the four DrugMatrix data sets each contained 842 compounds and had IRs of at least 16.5. Our model achieved its best predictive performance on the CHEMBL1909150 data set with an average ROC-AUC score of 0.978, indicating an outstanding ability to discriminate between the active and inactive classes. Furthermore, our model outperformed all GHOST models, which also demonstrated very high ROC-AUC scores on this data set. For CHEMBL1909157, our procedure obtained excellent results, with an ROC-AUC score of 0.866. Some of the GHOST models also achieved ROC-AUC scores greater than 0.8, but the GHOST model using GB performed significantly worse. Our model similarly outperformed all four GHOST comparison models on the CHEMBL1909132 and CHEMBL1909134 data sets.
@@
Figure 7 shows our model’s results on two highly imbalanced DS1 data sets, again compared with the GHOST algorithm paired with LR, RF, XGB, and GB. Recall from Section 3.1 that all DS1 data sets have an IR of 20.0 by construction. For CHEMBL100, our method achieved an outstanding ROC-AUC score of 0.974, while the highest-performing GHOST model (using LR) earned an ROC-AUC score of 0.959. While our BTDT-MBO method did not achieve the highest ROC-AUC score of comparison algorithms on CHEMBL8, it obtained the second-best ROC-AUC score, performing almost as well as the top-performing GHOST model using LR. All models in this case demonstrated ROC-AUC scores of at least 0.9, with the highest method earning a score of 0.979 and BTDT-MBO scoring 0.978.
@@
As outlined in Section 2.2, in addition to the default Gaussian kernel used in our BTDT-MBO method, we also conducted experiments that instead used the distance correlation to compute the weights between data elements. In our experiments, the BTDO-MBO method using a Gaussian weight function yielded higher ROC-AUC scores than the BTDT-MBO method using distance correlation for five out of the six data sets included in this work, with the BTDT-MBO algorithm using distance correlation achieving the best results on the CHEMBL1909157 DrugMatrix data set. However, the BTDT-MBO algorithm using a distance correlation weight function performed similarly to (and frequently almost as well as) the BTDT-MBO method using a Gaussian kernel on all six data sets, particularly on the DS1 data sets used for benchmarking. For the CHEMBL100 DS1 data set, the BTDT-MBO model using distance correlation yielded an ROC-AUC score of 0.970, slightly lower than BTDT-MBO using Gaussian weights (ROC-AUC = 0.974) but still higher than all comparison methods. Similarly, for the CHEMBL8 DS1 data set, BTDT-MBO using distance correlation scored 0.973, which was again lower than BTDT-MBO using a Gaussian weight function (ROC-AUC = 0.978) but higher than two comparison methods. All tested models scored very highly on this data set.
@@
Our model using distance correlation saw similar results on the DrugMatrix data sets. In fact, as mentioned above, BTDT-MBO using distance correlation achieved the best overall result of all tested models on the CHEMBL1909157 DrugMatrix data set, earning an ROC-AUC score of 0.878. BTDT-MBO using a Gaussian kernel scored an ROC-AUC value of 0.866; thus, both BTDT-MBO methods beat all tested GHOST models. For the CHEMBL1909150 data set, BTDT-MBO using distance correlation scored an ROC-AUC value of 0.975, higher than all comparison methods and only lower than our BTDT-MBO model using a Gaussian, which scored 0.978. On the CHEMBL1909132 DrugMatrix data set, BTDT-MBO using distance correlation earned an ROC-AUC score of 0.743, which places it below BTDT-MBO using Gaussian weights (ROC-AUC = 0.785) and GHOST using LR, but above the three other GHOST comparison methods. Our model using distance correlation saw similar relative results for the CHEMBL1909134 DrugMatrix data set, with an average ROC-AUC score of 0.766. The model earned the third-highest score of all methods tested for this data set, lower than our BTDT-MBO method using Gaussian weights (ROC-AUC = 0.792) and very close to GHOST paired with LR, which scored a 0.769. Overall, these results suggest that the distance correlation can be useful for quantifying similarity in molecular data sets, including those that are highly imbalanced. Furthermore, the distance correlation may be considered as a potential alternative weight function to the Gaussian function in similarity graph-based settings, particularly when the data is not close to a normal distribution.
To further visualize our proposed model’s performance on the benchmark data sets, we additionally utilized the residue-similarity (R-S) scores and associated R-S plots introduced by Hozumi et al. 47. For a given data point in a particular class, its residue score is defined as the sum of the distances from that point to points in other classes. The point’s similarity score is defined as the average distance from that point to points in its own class. In other words, the residue score measures how well data points in a given class are separated from other classes, and the similarity score measures how well data points in a given class are clustered together. Given a training/testing (or labeled/unlabeled) split of a data set, one can construct R-S score plots visualizing the elements of each class separately. In Figures 5 and 6, we plot R-S scores for a random labeled/unlabeled split of the six data sets used in our work. Points are separated into panels by their true class label, and the color of each point corresponds to its predicted class using our proposed BTDT-MBO model. We use the optimal weight function for each data set as discussed above.
Figure 5: R-S score plots for the four DrugMatrix data sets. The plots display R-S scores of unlabeled points from a random labeled/unlabeled partition of each data set. The $x$ and $y$ axes of the plots represent residue and similarity scores, respectively. From left to right, the panels plot points in class 1 (i.e., inactive compounds) and class 2 (i.e., active compounds). Each point is colored based on its class predicted by the proposed BTDT-MBO model.
Figure 6: R-S score plots for the two DS1 data sets. The plots display R-S scores of unlabeled points from a random labeled/unlabeled partition of each data set. The $x$ and $y$ axes of the plots represent residue and similarity scores, respectively. From left to right, the panels plot points in class 1 (i.e., inactive compounds) and class 2 (i.e., active compounds). Each point is colored based on its class predicted by the proposed BTDT-MBO model.
@@
Another novel addition to the proposed BTDT-MBO method is its threshold optimization step, the details of which are given in Section 2.1.1. This step chose the decision threshold that yielded the highest ROC-AUC score for a given partition of a particular data set. Because the partitions were created randomly, the class distribution of the unlabeled and labeled portions of a data set could vary. Additionally, due to the high level of class imbalance in the data, there may be very few data points from the minority class in the labeled or unlabeled set for a particular partition. Therefore, it is not surprising that in our experiments, the optimal thresholds for the splits varied considerably. However, the optimal thresholds often tended to be lower than the typical threshold of 0.5. For example, in one experiment on the CHEMBL1909157 data set (i.e., testing over 50 random labeled/unlabeled partitions), the lowest optimal threshold for the 50 random splits was 0.05, and the highest optimal threshold was 0.45. The median optimal threshold over all 50 splits was 0.15. These results reinforce the impact and the advantage of adjusting the decision threshold on classifier performance for highly imbalanced data.
3.4 Statistical Testing
In this paper, we perform statistical analysis using the experimental results; we perform the Friedman and the post-hoc Nemenyi test in order to achieve a more thorough evaluation. The results are shown in Table 6 and Figure 7, including the mean rank values, where smaller ranks demonstrate more competitive techniques. The critical distance was around 4.29, with the significance level being 0.05. Overall, the Friedman test’s null hypothesis is that all algorithms have the same performance. If the p-value is less that 0.05, the null hypothesis is rejected, which it was in our testing. Moreover, the Nemenyi post-hoc test is utilized to identify which procedures are significantly different from each other. Overall, it is shown that the proposed algorithm is ranked consistently the best among comparison methods.
3.5 Future Work and Limitations
While the proposed method has many advantages, such as the ability to perform accurately with highly imbalanced data, it has certain limitations similarly to any other algorithm. For example, the proposed procedure requires the calculation of a small number of the graph Laplacian’s eigenvectors and corresponding eigenvalues. This can obviously be computationally expensive for certain data sets. One possible direction in such a scenario is to construct an approximation of the full graph using other techniques, e.g., sampling-based procedures, in particular the Nyström Extension technique, detailed in sources such as 61, 62, 63. Such a technique can be utilized to very efficiently compute approximations of the eigenvalues and eigenvectors for the proposed procedure even for large data sets.
Future work regarding this study includes exploring imposing class size constraints into the technique. Moreover, we would like to explore integrating other correlation techniques, as well as incorporating extended-connectivity fingerprints or autoencoders into the algorithm structure.
Table 6:    Results for Statistical Testing
(a)    Reference for Tables 5b-5c
Method Method Method
‘1’ BTDT-MBO (Proposed) ‘4’ XGB-GHOST ‘7’ BT-SVM
‘2’ LR-GHOST ‘5’ BT-MBO ‘8’ BT-GB
‘3’ RF-GHOST ‘6’ GB-GHOST ‘9’ BT-RF
(b)    Results of Friedman’s Test
Chi-sq statistic p-value
112.39130 2.4237e-07
(c)    Mean rank values of the Friedman/Nemenyi test
‘1’ ‘2’ ‘3’ ‘4’ ‘5’ ‘6’ ‘7’ ‘8’ ‘9’
1.16667 2.16667 3,16667 3.66667 4.83333 6.33333 6.66667 8.00000 9.00000
Figure 7: The average rank diagram of the Friedman/Nemenyi test